OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 3i @triciaal

OpenStudy (triciaal):

you are given most of the zeros then find what else do you need to have real coefficients get rid of the complex number remember i^2 = -1 post what you get remember I haven't had time to try it yet. do you have options for the solution?

OpenStudy (anonymous):

f(x) = x4 - 6x3 + 12x2 - 90x + 416 f(x) = x4 - 45x2 + 180x - 416 f(x) = x4 - 35x2 + 180x - 416 f(x) = x4 - 6x3 - 12x2 + 90x - 416 are my answer choices @triciaal

OpenStudy (anonymous):

would you kind of solve it like the last one? make it like (x-4) (x+8) and idk how to do the other one

OpenStudy (triciaal):

yes

OpenStudy (triciaal):

on the phone give me a few minutes to be back

OpenStudy (triciaal):

(x-4)(x +8)(x -2 + 3 i)(x + 2-3 i)

OpenStudy (triciaal):

@jim_thompson5910 will you please help me with this?